If transactional money demand is mt= 0.4Y and speculative demand mz= 125 - 200r. What is the total money or how do you use these to calculate the LM equation?


1
Expert's answer
2021-07-30T05:31:02-0400

Solution


Md= 0.4Y

Ms= 125-200r

let money supply M=200


At equilibrium

Money supply= money demand


"200= 0.4Y+125-200r"


"0.4Y=200-125-200r"


The LM equation is:


"Y=187.5+500r"